OpenStudy (roadjester):

Find the arc length function for the curve with starting point (0,1).

OpenStudy (blockcolder):

What curve?

OpenStudy (roadjester):

\[y=\sin^{-1} x + \sqrt{1-x^2}\]

OpenStudy (blockcolder):

\[s(x)=\int_{0}^{x} \sqrt{1+\left ( \frac{1}{\sqrt{1-x^2}}-\frac{x}{\sqrt{1-x^2}}\right )^2}\ dx \\ =\int_{0}^{x} \sqrt{1+\left ( \frac{1-x}{\sqrt{1-x^2}}\right )^2}\ dx\\ =\int_{0}^{x} \sqrt{1+\frac{(1-x)^2}{1-x^2}}\ dx\\ =\int_0^{x} \sqrt{1+\frac{1-x}{1+x}}\ dx\\ =\int_0^{x} \sqrt{\frac{2}{1+x}}\ dx\] to be continued....

OpenStudy (blockcolder):

\[s(x)=\sqrt{2}\int_0^x (1+x)^{-1/2}\ dx\\ =2\sqrt{2} \sqrt{1+x}|_0^x\\ =2\sqrt2(\sqrt{1+x}-1)\] Whew!